12/5x+7=31-3/5x
-2x+5x-9=3(x-4)-5
-5/6x=-15
6=x/7
4/5x=-15

Answer by rfer(16322)

12/5x+7=31-3/5x
15/5x=24
5x=1.6
x=0.32
==========